Please help with a step by step solution for homework problem for chapter 16 . The textbook I am using is Fundamental of Cost Accounting sixth edition. My question is in reference to exercise 16-65. Part A and Part B. The answer I submitted was incorrect

McDormand, Inc reported a $3,200 unfavorable price variance for variable overhead and a $32,000 unfavorable price variance for fixed overhead. The flexible budget had $1,080,000 variable overhead based on 36,000 direct labor hours; only 34,080 were worked. Total actual overhead was $1,829,600. The number of estimated hours for computing the fixed overhead application rate totaled 38,600.

Required:

a. Prepare a variable overhead price analysis :

   Price variance:

   Efficiency variance:

   Variable overhead cost variance:

b. Prepare a fixed overhead analysis:

   Price variance:

   Production volume variance:

   Fixed overhead cost variance:

Your help is very much appreciated.
Answer #1

Solution:

A)

1) Variable overhead price variance = (Actual hours * Standard variable overhead rate ) - Actual variable overhead cost

(-)3,200 = (34,080 *1,080,000/36,000) -Actual variable overhead cost

Actual variable overhead cost = 1,022,400 +3,200

Actual variable overhead cost = 1,025,600

2) Variable overhead efficiency variance =(Standard hours for actual output - Actual hours ) + Standard variable overhead rate

=(36,000 - 34,080)*1,080,000/36,000

=1,920 * 30

=57,600(Favorable)

3) Variable overhead cost variance = (Standard hours for actual output * Standard variable overhead rate) - Actual variable overhead

=(36,000 * 1,080,000/36,000 ) - 1,025,600

=1,080,000 - 1,025,600

=54,400(favorable)

Variable overhead price variance 3,200 (unfavorable)
Variable overhead efficiency variance 57,600(favorable)
variable overhead cost variance 54,400(Favorable)

B)

1)Fixed overhead price variance = Budgeted fixed overhead - Actual fixed overhead cost

(-)32,000 = Budgeted fixed overhead -(1,829,600 - 1,025,600)

Budgeted fixed overhead = 804,000 - 32,000

Budgeted fixed overhead = $ 772,000

Standard fixed overhead rate = 772,000 / 38,600

= 20

2) Fixed overhead volume variance = (Standard hours for actual output - Budgeted hours)*Standard fixed overhead rate

=(36,00 - 38,600) *20

= 52,000(Unfavorable)

3)Fixed overhead cost variance = Fixed overhead price variance + Fixed overhead volume variance

=(-)32,000+(-)52,000

=84,000

Fixed overhead price variance $32,000(Unfavorable)
Fixed overhead volume variance $52,000(Unfavroable)
Fixed overhead cost variance $84,000(unfavorable)